# Tightly Coupled SLAM with Imprecise Architectural Plans

### v2

@@ VI Graph Optimization
Before the S-Graph and A-Graph are coupled, the S-Graph state at time can be defined as:
Similarly, the A-Graph state is defined as:
where are the robot poses at selected keyframes in the S-Graph frame of reference, are the plane coefficients of the and wall-surfaces of the S-Graph and A-Graph respectively, contains the poses of the and four-wall rooms of the S-Graph and A-Graph respectively. are the poses of the two-wall rooms in the S-Graph. contains the poses of doorways of the A-Graph, are the floors levels, and models the drift between the odometry frame and the S-Graph reference frame . If at time there is no match obtained between the A-Graph and S-Graph we perform single S-Graph optimization as detailed in [7].
@@
Transformation Estimation: Upon receiving the match (Section IV) and performing graph coupling (Section V), we combine Eq. 4 and Eq. 5 to make a global state with additional transformation factor . represents the transformation between the origins of the A-Graph and the S-Graph. It is important to note that at this stage, the deviated wall-surface and room are not included for estimating .
Deviation Estimation: After optimizing we already have an initial guess of the transformation between the A-Graph and the S-Graph, and we can incorporate the deviated wall-surface and room entities into the graph with appropriate deviation factors. Our state then becomes where are the deviation factors between wall-surfaces and are the deviation factors between rooms. When optimizing we keep constant to obtain a good initial estimation of the deviation between the matched deviated entities.
@@ VII-A Methodology
Setup. We evaluated our algorithm in both simulated and real environments. Both simulated and real experiments were performed using a laptop computer with an Intel i9-11950H (8 cores, 2.6 GHz) with 32 GB of RAM.
Baselines. We have selected the following LiDAR-based baselines for comparisons due to their suitability, their reported results, and the availability of their code: AMCL [9], UKFL [13], OGM2PGM [12], IR-MCL [14], and iS-Graphs [8]. Although some works do localization in imprecise architectural plans, the absence of open-source implementations prevents direct comparison with these methods. As each selected baseline takes a different map input for localization, we used 2D occupancy grid maps for AMCL and OGM2PGM, 3D meshes for UKFL and IR-MCL, and A-Graphs for iS-Graphs and our method diS-Graphs, respectively.
Simulated Datasets. We validate the algorithms in five simulated datasets named SE1 to SE5. To record the datasets, we use the Gazebo physics simulator to recreate the robot, its sensors (LiDAR), and the 3D indoor environments obtained from actual architectural plans. We report the absolute trajectory error (ATE) compared with the available ground truth trajectory. We also report the localization convergence success rate of all methods.
Real Datasets. We collected data with a legged robot equipped with a Velodyne VLP-16 LiDAR, at five different construction sites (RE1 to RE5), with existing architectural plans. In real experiments, we report the Root Mean Square Error (RMSE) between estimated 3D maps and ground truth architectural plans due to the absence of pose ground truth. The RMSE, while a mapping metric, primarily demonstrates the localization accuracy of all baselines, as mapping accuracy inherently reflects pose estimation quality. Furthermore, we report all methods’ convergence rates and convergence and computation times.
Deviations. Note that, given the absence of actual ground truth deviations in a real environment with respect to the plans, we explicitly deviate the wall-surfaces entities in the architectural plans to have a ground truth estimate of the deviations for both simulated and real datasets. For all datasets, we performed five separate tests introducing uniform random wall-surface deviations in architectural plans ranging from to in translation, and and in rotations. Moreover, to test the robustness of our algorithm against deviations, we conducted 25 experiments per dataset with varying deviations.
Ablation. We conducted ablation studies on two key components of our algorithm. 1. Covariance assignment: As mentioned in the graph matching and graph merging (Section. IV & V), we assign covariances to the deviation factors of the semantic elements, based on their deviation likelihood. Here, we analyze the effect of assigning equal covariances, referred to as uniform covariances (UC), to both deviated and non-deviated elements. 2. Alternating optimization: In graph optimization (Section. VI) we discuss the use of alternating optimization for simultaneous localization and deviation estimation. Here, we study the performance when using only a single optimization cycle (SO).
@@ VII-B Results and Discussion
Bold values are the best and the second best are underlined. ‘-’ refers to an unsuccessful run.
@@
Fig. 6(a) summarizes the ATE performance of all the baseline algorithms. AMCL shows the highest median ATE and a relatively narrow distribution, indicating consistently high error rates. UKFL and OGM2PGM demonstrate moderate performance with similar median ATEs, although OGM2PGM shows a wider range of errors. IR-MCL exhibits the largest variability, suggesting inconsistent performance in different scenarios. AMCL, OGM2PGM, IR-MCL, UKFL and iS-Graphs assume no deviations between “as-planned” and “as-built” environments. These methods attempt to match laser scans with an incorrect reference map, leading to consistent misalignment. Therefore, they show higher ATEs across datasets, indicating their vulnerability to architectural deviations. diS-Graphs can detect and estimate the deviations between “as-planned” and “as-built” environments, resulting in lower ATE than other baselines.
Point Cloud Alignment Error. Table II shows the RMSE of the point clouds with respect to the ground truth for all methods and ours. In case of deviations in construction from the plans, diS-Graphs shows better accuracy than AMCL, better than OGM2PGM, better than IR-MCL, and better than iS-Graphs. Although UKFL’s average error is equal to diS-Graphs’, it has a very low convergence rate (see Table III) rendering the comparison unfair. Moreover, it cannot estimate the deviations between “as-planned” and “as-built” environments. Fig. 6(b) summarizes the performance of all algorithms in real environments. AMCL exhibits the highest median and widest interquartile range, indicating greater variability in performance. IR-MCL presents a large spread of results, while OGM2PGM shows moderate performance with a smaller range of variability compared to AMCL and IR-MCL. Although UKFL and diS-Graphs show the lowest median RMSE, suggesting superior accuracy, the comparison with UKFL is not fair because of its low convergence rate. Because of our simultaneous estimation of deviations and initial transformation, we can not only simultaneously globally localize the robot but also estimate the deviations between semantic elements of “as-planned” and “as-built” environments moreover improving the overall accuracy compared to other algorithms.
@@
Convergence and Computation Time. Table IV shows the convergence time for each algorithm. The convergence time is the time it takes the algorithm to globally localize. IR-MCL and OGM2PGM have the best convergence time. iS-Graphs and diS-Graphs have considerably longer convergence times because they need to detect a certain number of rooms in the environment for graph matching to find a unique match. In addition, the previous version of graph matching struggled to resolve symmetries during the matching process, resulting in longer convergence times. However, the modifications in the graph matching algorithm (Section IV) proposed in this work improve the symmetry resolution ability and reduce the convergence time by almost .
Table IV shows the computation time for each algorithm. The computation time is the time used for each pose update. On average, the computation time of our algorithm is milliseconds, showing its real-time performance. Our computation time is the best compared to other 3D algorithms. OGM2PGM and AMCL have the best computation time because, unlike others, they process 2D information.
Deviation Estimation. Fig. 8 shows the amount of deviation our algorithm can correctly estimate in real environments. The minimum detectable deviation is bounded by our LiDAR sensor’s accuracy. With our LiDAR having an accuracy of [27] and considering that of measurements fall within three standard deviations () of the mean, our system can reliably detect deviations only when they exceed . The maximum translational and rotational deviation in wall-surfaces our algorithm can detect accurately is and respectively. Fig. 5 shows several qualitative results from our diS-Graphs method for real experiments. The robot can successfully localize, map, and estimate deviations in these environments. Note that we only show the rooms and walls for better understanding, and all the other semantic elements and the robot are not shown. We do not show RE5 in Fig. 5, as the robot could not localize itself in this sequence.
@@ VIII Conclusions and Future Work
This paper presents a novel approach to tightly couple SLAM with imprecise architectural plans. Our algorithm establishes direct correspondences between structural elements in “as-built” environments and their “as-planned” counterparts, enabling simultaneous robot localization and deviation estimation between the two representations. By detecting and estimating the deviations between “as-built” and “as-planned” environments through tight coupling, our method outperforms the current best approach with 43% better localization in simulations, 7% improved mapping accuracy in real environments, and enhanced robustness to architectural deviations. Additionally, our algorithm provides an estimate of existing deviations up to in translation and in rotation. Our algorithm is limited by the need to have enough distinctive semantic elements (i.e. wall-surfaces and rooms) to provide a unique match. As future work, we plan to add the ability to detect and match more semantic elements, which will translate into an improvement in the convergence rate and deviation detection range. Moreover, to gain flexibility, we plan to improve the matching process by adding the ability to detect rooms consisting of more than four walls.

### v3

@@ VI Graph Optimization
Before the S-Graph and A-Graph are coupled, the S-Graph state at time can be defined as [7]:
Similarly, the A-Graph state is defined as [8]:
where models the drift between the odometry frame and the S-Graph reference frame , and the other variables are described in Section III. Note, if at time there is no match between the A-Graph and S-Graph, we perform single S-Graph optimization as detailed in [7] until we detect it.
@@
Transformation Estimation: Upon receiving the match (Section IV) and performing graph coupling (Section V), we combine Eq. 4 and Eq. 5 to make a global state with additional transformation . It is important to note that at this stage, the deviated wall-surface and room are not included for estimating .
Deviation Estimation: After optimizing we already have an initial guess of the transformation between the A-Graph and the S-Graph, and we can incorporate the deviated wall-surface and room entities into the graph with appropriate deviation factors. Our state then becomes When optimizing we keep constant to obtain a good initial estimation of the deviation between the matched deviated entities.
@@ VII-A Methodology
Setup. We evaluated our algorithm in both simulated and real environments. Both types of experiments were performed using a laptop computer with an Intel i9-11950H (8 cores, 2.6 GHz) with 32 GB of RAM.
Baselines. We have selected the following LiDAR-based baselines for comparisons due to their suitability, their reported results, and the availability of their code: AMCL [9], UKFL [13], OGM2PGM [12], IR-MCL [14], and iS-Graphs [8]. Although other works do localization in imprecise architectural plans, the absence of open-source implementations prevents direct comparison with these methods. As each selected baseline takes a different map input for localization, we used 2D occupancy grid maps for AMCL and OGM2PGM, 3D meshes for UKFL and IR-MCL, and A-Graphs for iS-Graphs and our method diS-Graphs, respectively.
Simulated Datasets. We validate the algorithms on five simulated datasets (SE1 to SE5) using Gazebo to recreate the robot, LiDAR sensor, and 3D environments from actual architectural plans. We report absolute trajectory error (ATE) and localization convergence success rates.
Real Datasets. We collected data using a legged robot with Velodyne VLP-16 LiDAR at five construction sites (RE1 to RE5) with existing architectural plans. We report Root Mean Square Error (RMSE) between estimated 3D maps and ground truth plans (indicating localization accuracy since mapping quality reflects pose estimation), plus convergence rates and computation times.
Deviations. Note that, given the absence of actual ground truth deviations in a real environment with respect to the plans, we explicitly deviate the wall-surfaces in the architectural plans to have a ground truth estimate of the deviations for both simulated and real datasets. For all datasets, we performed five separate tests introducing uniform random wall-surface deviations in architectural plans ranging from to in translation, and and in rotations. Moreover, to test the robustness of our algorithm against deviations, we conducted 25 experiments per dataset with varying deviations.
Ablations. We conducted ablation studies on two key components of our algorithm. 1. Covariance assignment: As mentioned in the graph matching and graph merging (Section. IV & V), we assign covariances to the deviation factors of the semantic elements, based on their deviation likelihood. Here, we analyze the effect of assigning equal covariances, referred to as uniform covariances (UC), to both deviated and non-deviated elements. 2. Alternating optimization: In graph optimization (Section. VI) we discuss the use of alternating optimization for simultaneous localization and deviation estimation. Here, we study the performance when using only a single optimization cycle (SO).
@@ VII-B Results and Discussion
Fig. 6(a) summarizes the ATE performance of all the baseline algorithms. AMCL shows the highest median ATE and a relatively narrow distribution, indicating consistently high error rates. UKFL and OGM2PGM demonstrate moderate performance with similar median ATEs, although OGM2PGM shows a wider range of errors. IR-MCL exhibits the largest variability, suggesting inconsistent performance in different scenarios. AMCL, OGM2PGM, IR-MCL, UKFL and iS-Graphs assume no deviations between “as-planned” and “as-built” environments. These methods attempt to match laser scans with an incorrect reference map, leading to consistent misalignment. Therefore, they show higher ATEs across datasets, indicating their vulnerability to architectural deviations. Our diS-Graphs can detect and estimate the deviations between “as-planned” and “as-built” environments, resulting in the lowest ATE.
Point Cloud Alignment Error. Table II shows the RMSE of the point clouds with respect to the ground truth for all baseline methods and ours. In the presence of deviations between “as-planned” and “as-built”, diS-Graphs shows better accuracy than AMCL, better than OGM2PGM, better than IR-MCL, and better than iS-Graphs. Although UKFL’s average error is equal to diS-Graphs’, it has a very low convergence rate (see Table III) rendering the comparison unfair. Moreover, it cannot estimate the deviations between “as-planned” and “as-built” environments. Fig. 6(b) summarizes the performance of all algorithms in real environments. AMCL exhibits the highest median and widest interquartile range, indicating greater variability in performance. IR-MCL presents a large spread of results, while OGM2PGM shows moderate performance with a smaller range of variability compared to AMCL and IR-MCL. Although UKFL and diS-Graphs show the lowest median RMSE, suggesting superior accuracy, the comparison with UKFL is not fair because of its low convergence rate. Because of our simultaneous estimation of deviations and initial transformation, we can not only simultaneously globally localize the robot but also estimate the deviations between semantic elements of “as-planned” and “as-built” environments, moreover improving the overall accuracy compared to other algorithms.
@@
Convergence and Computation Time. Table IV shows the convergence time for each algorithm. The convergence time is the time it takes the algorithm to globally localize. IR-MCL and OGM2PGM have the best convergence time. iS-Graphs and diS-Graphs have considerably longer convergence times because they need to detect a certain number of rooms in the environment for graph matching to find a unique match. The graph matching of [8] struggles to resolve symmetries during the matching process, resulting in longer convergence times. However, the modifications in the graph matching algorithm (Section IV) proposed in this work improve the symmetry resolution ability and reduce the convergence time by almost . Table IV shows the computation time for each algorithm. We specifically report the time spent for each pose update. On average, the computation time of our algorithm is milliseconds per pose update, showing real-time performance. Our computation time is the best compared to other 3D algorithms. OGM2PGM and AMCL. The smaller computation time of the others is due to their simpler optimization in 2D.
Deviation Estimation. Fig. 8 shows the amount of deviation our diS-Graphs can correctly estimate in real environments. The minimum detectable deviation is bounded by the LiDAR resolution. With our LiDAR errors having a stamdard deviation of [27], and considering that of measurements fall within three standard deviations, our diS-Graph can reliably detect deviations only when they exceed . The maximum translational and rotational wall-surface deviation that our diS-Graphs can detect is and respectively. Fig. 5 shows several qualitative results from our real experiments. The robot can successfully localize, map, and estimate deviations in these environments. Note that we only show the rooms and walls for better understanding, and all the other semantic elements and the robot are not shown. We do not show RE5 in Fig. 5, as the robot could not localize itself in this sequence.
@@ VIII Conclusions and Future Work
This paper presents a novel approach to tightly couple SLAM with imprecise architectural plans. Our algorithm establishes direct correspondences between structural elements in “as-built” environments and their “as-planned” counterparts, enabling simultaneous robot localization and deviation estimation between the two representations. By detecting and estimating the deviations between “as-built” and “as-planned” environments through tight coupling, our method outperforms the current best approach with 43% better localization in simulations, 7% improved mapping accuracy in real environments, and enhanced robustness to architectural deviations. Our algorithm provides an estimate of existing deviations up to in translation and in rotation. Our algorithm is limited by the need for sufficient distinctive semantic elements (i.e., wall-surfaces and rooms) to provide unique matches. As future work, we plan to detect and match more semantic elements and support rooms with non-rectangular geometries, including curved walls and partial floorplans. Additionally, we aim to integrate multi-modal sensing (RGB-D, vision-based semantic segmentation) to enhance robustness in visually degraded environments. These improvements will enhance the convergence rate and enable deviation detection for additional detected entities beyond just rooms and walls.